# Depth-Resolved Characterization of Centrifugal Disk Finishing of Additively Manufactured Inconel 718

### v1

@@ 4 Discussion
Nonetheless, the results described in section 3.1 suggest avenues for process optimization for achieving an ideal surface texture within a short duration. The 50:50 bimodal mixture of media was able to provide an appreciable improvement in measures of surface roughness, e.g., $R_{a}=$ 1.15 $\mu$m after 75 mins of processing compared with the starting value of $R_{a}=$ 4.9 $\mu$m before finishing. This measure would be eventually superseded by the smaller 6$\times$6 SCT media that provided $R_{a}=$ 0.79 $\mu$m after 135 mins of finishing. Utilizing the bimodal media alone would have provided $R_{a}=$ 0.97 $\mu$m after 135 mins of processing. This suggests that dynamically altering the media from bimodal to unimodal may facilitate an adequate amount of processing in a shorter amount of time. Nonetheless, such studies do not provide a comparison of the surface resulting from such processing routes beyond a trivial comparison of its surface roughness measures like $R_{a}$, $R_{q}$, and $R_{z}$. This detail has been realized to be critical for formulating predictive models of part functional response [10, 17]. To bridge this knowledge gap, a detailed analysis of the surfaces resulting after the equivalent processing times but with different media is attempted in the next section.
@@ 4.1 Fourier Analysis
Figure 9 shows representative spectrums of the NF (Fig. 9a ), U1 (Fig. 9b), U2 (Fig. 9c), B1 (Fig. 9d), and B2 (Fig. 9e) surfaces with insets showing the low wavenumber regions in more detail. In the figure, we see the NF specimen surface has a high amount of disorder, evidenced from the high amplitudes of waves with low wavenumbers (inset in Fig. 9a). This disorder originates from the layer-by-layer nature of additive manufacturing which produces semi-regular undulations on the surface. In addition, partially melted powders also fuse to the surface during AM, which can be an additional source of the disorder in this surface. The quantification of spectral entropy of this surface suggested $H=0.43$. Upon finishing with the 6x6 SCT media for 75 mins. (U1), the specimen surface assumed a less disordered state evidenced in its much lower spectral entropy of H=0.27 (Fig. 9b). However, upon further finishing with the same media to a total of 135 mins, the spectral entropy of the surface increased to H=0.47 (U2). In comparison, finishing with the bimodal media resulted in increase in spectral entropy to H = 0.48 (B1, 75 mins), and H = 0.52 (B2, 135 mins).
The results summarized in the previous paragraph suggest that the surfaces resulting from CDF with bimodal media were always more disordered than surfaces resulting from finishing with unimodal media. This disorder can originate from multiple sources. The first phase of finishing by the 6$\times$6 media for 75 mins. (U1) reduces the disorder compared with the NF surface, in-line with the also observed reduction in surface roughness measures. However, the additional 60 mins of finishing with the same 6$\times$6 media (U2) results in a further decrease in surface roughness and an increase in the spectral entropy. This implies that the mechanics of surface smoothing underwent a transition between 75, and 135 mins, which resulted in this increase of disorder. This transition was accompanied by an increase in the magnitude of residual stresses at $\sim$70 $\mu$m for the U2 specimen. A similar transition was however not seen in the bimodal media that exhibited a decrease in surface roughness and increase in spectral entropy throughout the finishing process.
@@
5 Summary
A unique combination of traditional surface profilometry and synchrotron high-energy X-ray diffraction were combined to build a more complete picture of surface roughness reductions and changes to microstructure below the surface in AM Incoel 718. Four different centrifugal disk finishing conditions (two media types $\times$ two finishing times) were examined along with a reference specimen. From the characterization, we found:
@@ 5 Summary
    Both abrasive media types removed strongly (fiber) textured material at an AM surface, leaving only bulk (cube) textures without the introduction of any texture components.

### v2

@@ 4 Discussion
The finishing mechanism of CDF involves a combination of shear-based material removal and smearing of surface roughness asperities through plastic deformation, both of which result in smaller measures of surface roughness. These mechanisms involve unique trajectories of microstructure evolution that produce different microstructural consequences. Furthermore, the exact mechanism is dynamically governed by the local mechanical conditions during CDF. Here the results indicate that the amount of plastic deformation induced by CDF is moderate (at least for the processing conditions examined). The increase in diffraction peak widths observed near the surface is tied to an increase in local dislocation content and deformation compatibility which occurs during plastic deformation. However, the amount of plastic deformation was not sufficient to induce large-scale grain rotation or dynamic recrystallization which would lead to the appearance of new crystallographic texture components in the pole figures in comparison to the not-finished condition. These observations are in contrast to most shear-based material removal processes which are usually classified as surface severe plastic deformation processes (like machining and burnishing). Recent efforts have attempted to phenomenologically model these phenomena by investigating the coupled evolution of part surface roughness and the media condition with respect to CDF processing time [12]. As the removed surface material here had a relatively strong fiber texture, which will naturally have a strongly anisotropic mechanical response, the effects of its removal will be load-condition dependent. The increased defect content and accompanying work-hardening at the surface will produce a moderate ‘case-hardening’ effect that can be beneficial in high-cycle fatigue applications.
The results described in section 3.1 suggest avenues for process optimization for achieving an ideal surface texture within a short duration. The 50:50 bimodal mixture of media was able to provide an appreciable improvement in measures of surface roughness, e.g., $R_{a}=$ 1.15 $\mu$m after 75 mins of processing compared with the starting value of $R_{a}=$ 4.9 $\mu$m before finishing. This measure would be eventually superseded by the smaller 6$\times$6 SCT media that provided $R_{a}=$ 0.79 $\mu$m after 135 mins of finishing. Utilizing the bimodal media alone would have provided $R_{a}=$ 0.97 $\mu$m after 135 mins of processing. This suggests that dynamically altering the media from bimodal to unimodal may facilitate an adequate amount of processing in a shorter amount of time. Nonetheless, such studies do not provide a comparison of the surface resulting from such processing routes beyond a trivial comparison of its surface roughness measures like $R_{a}$, $R_{q}$, and $R_{z}$. This detail has been realized to be critical for formulating predictive models of part functional response [13, 28]. To bridge this knowledge gap, a detailed analysis of the surfaces resulting after the equivalent processing times but with different media is attempted in the next section.
@@ 4.1 Fourier Analysis
Figure 11 shows representative spectrums of the NF (Fig. 11a ), U1 (Fig. 11b), U2 (Fig. 11c), B1 (Fig. 11d), and B2 (Fig. 11e) surfaces with insets showing the low wavenumber regions in more detail. In the figure, we see the NF specimen surface has a high amount of disorder, evidenced from the high amplitudes of waves with low wavenumbers (inset in Fig. 11a). This disorder originates from the layer-by-layer nature of additive manufacturing which produces semi-regular undulations on the surface. In addition, partially melted powders also fuse to the surface during AM, which can be an additional source of the disorder in this surface. The quantification of spectral entropy of this surface suggested $H=0.43$. Upon finishing with the 6$\times$6 SCT media for 75 mins, (U1), the specimen surface assumed a less disordered state evidenced in its much lower spectral entropy of H=0.27 (Fig. 11b). However, upon further finishing with the same media to a total of 135 mins, the spectral entropy of the surface increased to H=0.47 (U2). In comparison, finishing with the bimodal media resulted in increase in spectral entropy to H = 0.48 (B1, 75 mins), and H = 0.52 (B2, 135 mins).
The results summarized in the previous paragraph suggest that the surfaces resulting from CDF with bimodal media were always more disordered than surfaces resulting from finishing with unimodal media. This disorder can originate from multiple sources. The first phase of finishing by the 6$\times$6 media for 75 mins, (U1) reduces the disorder compared with the NF surface, in-line with the also observed reduction in surface roughness measures. However, the additional 60 mins of finishing with the same 6$\times$6 media (U2) results in a further decrease in surface roughness and an increase in the spectral entropy. This implies that the mechanics of surface smoothing underwent a transition between 75, and 135 mins, which resulted in this increase of disorder. This transition was accompanied by an increase in the magnitude of residual stresses at $\sim$70 $\mu$m for the U2 specimen. A similar transition was however not seen in the bimodal media that exhibited a decrease in surface roughness and increase in spectral entropy throughout the finishing process.
@@
5 Summary and Conclusions
A unique combination of traditional surface profilometry and synchrotron high-energy X-ray diffraction were combined to build a more complete picture of surface roughness reductions and changes to microstructure below the surface in AM Inconel 718. Four different centrifugal disk finishing conditions (two media types $\times$ two finishing times) were examined along with a reference specimen. From the characterization, we found:
@@ 5 Summary and Conclusions
    Both abrasive media types removed strongly (fiber) textured material at an AM surface, leaving only bulk (cube) textures without the introduction of any new crystallographic texture components.
@@
The results here point to the need for future work to understand the effect of preexisting geometries on the mechanics of CDF and their effects of fatigue life. Complicated geometries will significantly influence the mechanics of flow of the media during CDF, and thereby alter its process consequences. Without a thorough understanding of these underpinnings, it is not clear how process optimization can be achieved for parts with complex geometries, such as those commonly originating from additive manufacturing. In addition, CDF appears to provide an interesting combination of material removal with only moderate plastic deformation. How this combination affects fatigue life should be further explored.